RC Update and Run plan for Sep 20,2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are M-F at 4:15 PM

To join the Meeting:  https://bluejeans.com/861439920
To join via phone :  +1.408.740.7256 (US (San Jose)), +1.888.240.2560 (US Toll Free))
** This page is in progress. **

Sunday 20 Sept 2020.

General updates

(See d2n Run Plan 2020 for broader status/overview.)

  • d2n program completed around 9:30am (ran about 1 shift longer than initially planned)
    • After looking at the He3 elastics data, it was decided to run 8 hour longer with some more aggressive paddle settings to improve statistics [1]
    • E:QE ratio improved [2]; we'll have to work out the final numbers offline.
    • Target polarization continues to improve (good!), NMR calibration constant is likely overly optimistic, needs updating (but we'll stick with it for now to maintain consistent relative numbers)
  • Had an issue with s1y05+(B) signals were missing after paddles turned back on in the morning
    • HV readbacks seemed to be good, thought this might be a issue with CAEN firmware, or in the DAQ/software chain. No problems found with signals when I got in to check -- puzzling.
    • Simona reports this is a known issue (tube needs to 'warm up' or something) and should call her. Strange symptom... [3]
  • Polarized 3He Elastic Longitudinal Asymmetry Study begins (E12-06-121A)
    • 9:00am Rotated HMS to 17deg, and checked analysis cuts and paddle settings
    • 9:50am Rotated SHMS to 13deg, no opportunity to check settings yet; beam went off
  • 9:45am Beam off for MCC/Cryo to investigate 1L20 turbo cube issue (insulating vacuum) [4], and trim card problem [5]
    • 1L20 Turbo Repaired/Reset by 10:30am
    • MQA8A20 trim card issue identified later in this down
  • 11:30 Beam back --- Production for E12-06-121A continues
  • 5:00pm Found lasers had tripped in day shift; apparently around 10:15am [6]
    • Noticed at next NMR at 4:55pm showed low polarization (~30%) [7]
    • Took escorted access to inspect interlocks on target panels [8] and found that a taped on target panel had shifted enough to disengage an interlock switch. (Same panel but different interlock switch(P5) than the other day) [9]
    • Added an alarm on the pumping chamber temp as a 'Laser Trip' alarm for now [10]
    • Had to go in to fix the same switch again at 11pm [11]
  • Elected to be aggressive on HMS paddle On/off config to improve E:QE ratio, and allow us to push to 30uA to fill the pipe on the HMS; trading some potential complexity in the analysis for a big boost in the FoM [12]
  • A number Accelerator issues during OWL; but managed ~4 hours of production before program ended at 7am
  • Beam will shut off at 7am, Monday morning for 1 hour of Beam Study tests to address energy stability issues

Notes for the Future

  • Hall will move to Restricted Access on Monday
    • Expectation is 11am--noon; depends when RadCon is available
    • EPR photodiode will be replaced on Monday afternoon (JP, Junhao, RadCon)
  • Bill, Murchhana check compass/field measurements
    • Also checking with Jixie re: modeled fields
    • Need to make a write-up that clearly defines all the different coordinate systems used, swapped power supply/polarity issues, etc.

Pending Issues

  • Some beam energy stability issues have been noticed by Hall B (in particular); Jay is looking into this (Not immediate problem, just leaving this here as notes for future)
    • These should not impact Hall C significantly, but may feed into beam polarization corrections.
    • Jay has link some of this discussion to the Hall C beam energy measurement HCLog entry and cross posted to the HCLog. This will help us follow breadcrumbs during beam polarization analysis for 1-pass.
    • Hall C beam energy drifted up after the Hall C beam energy measurement was taken [13]
    • Hall B Moller Fri, 09/18/2020 - 10:20 was higher than expected: -86.233 +/- 1.494 (several percent) [14]
    • Energy shift that would impact Hall B polarization at few percent level [15]
    • Energy lock switched from 2A to 3C Sat (19th) @ 9:40 [16]. No notice to Halls, unfortunately..

Current Run Plan

Saturday DAY → Monday OWL

  • NOTE: Hall C is in control of the Charge Asymmetry feedback system until Monday
    • Charge Feedback Parity DAQ Instructions
    • We will have to run the ./FlipPCvaluesHallC script in the Parity DAQ VNC session after every IHWP change

Monday DAY (7am)

  • Rotate SHMS to 18.0 degrees and set to -5.6 GeV/c when MCC takes over (7am).
    • Monitor SOLID test stand on camera (as well as angle) during rotation.
    • Stay at longitudinal/0 for initial tests. JP will provide details to Target group.
  • Everything is opening up, so RadCon is pretty busy and isn't sure how quickly they will get to Hall C.
    • Best guess is they'll be able to open the Hall to Restricted around 11--noon and be able to support the EPR diode swap in the afternoon

Standing Orders

  • Change the IHWP at least every 8 hours when running on glass cells. Call Halls A and B in advance. If there are any issues, call the RC
    • NOTE: Hall C is in control of the Charge Asymmetry feedback system until Monday Charge Feedback Parity DAQ
    • We will have to run the ./FlipPCvaluesHallC script in the Parity DAQ VNC session after every IHWP change
  • Ensure that all information in the CODA run-start GUIs is correct for every run!
  • If beam goes away for more than 3h, or MCC does operations on the beam that may change the beam tune, then move target to No Target position,
    • verify Harps scans at 3H07A and 3H07B are within spec, and
    • ensure beam is stable (with all locks ON) at 0, 0mm before continuing.
    • Beam recovery procedure has more detailed procedure if needed.
  • New beam positions have been verified and established to be (0,0) mm for all targets.
    • Monitor that beam is at (0,0) using the stripchart on the Wall screen and verify no scraping on the glass cells using the Analysis monitor.
  • Check the position locks are on.
  • Check that CODA is running normally
  • Check online replays shortly after starting runs, & full replay results too.
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.
!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Follow the updated guidelines in the How-to#Spectrometer_Magnets Spectrometer Magnets How to
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • If any magnet trips, snapshot the associated power supply window and post it to hclog before resetting the interlock.
  • Monitor the beam position on target and call MCC immediately if you see large excursions (1 mm level).
  • When rotating the spectrometer, use HCCAM01 to keep an eye on the SOLID Cherenkov test stand when rotating SHMS (in addition to watching the live angle-cam image, etc). The updated rotation instructions are here: https://hallcweb.jlab.org/wiki/index.php/Spectrometer_Rotation#NOTICES_IN_EFFECT

Useful Links

PolHe3 Polarization Measurements

Misc Links